2. Josephine and Hilda are sisters. hilda is 17yrs. older than Josephine, In 3yrs.
Hildas age will be twice Josephine's age.
Q: what are their age in 3yrs. from Now?

Answer by ptaylor(2198)

Let x=Josephine's age in three years
Then x+17=Hilda's age in three years
Sooo
x+17=2x subtract x from each side
x=17----Josephine's age in three years
x+17=17+17=34---Hilda's age in three years
Hope this helps---ptaylor